OpenStudy (brittany):

rational or irrationl 5√64

OpenStudy (anonymous):

As it stands irrational. Simplified to 40, rational.

OpenStudy (anonymous):

\[5\sqrt{8*8}=5*8=40\] rational

OpenStudy (anonymous):

rational

OpenStudy (amistre64):

if you can get rid of that sqrt part, its rational; sqrt(64) = 8, so its rational